OpenStudy (anonymous):

Using the comparison test to determine if the following is convergent or divergent: Sum of [4(sqrt(n))-1]/(n^(5/4)+2n-1)

OpenStudy (aum):

\[ \frac{4\sqrt{n}-1}{n^{5/4}+2n-1} \]

OpenStudy (aum):

The largest power in the numerator is 1/2 and that in the denominator is 5/4. Therefore, this series will behave much like \[ \frac{\sqrt{n}}{n^{5/4}} = \frac{1}{n^{3/4}} \]which diverges as a p-series with p <= 1 diverges.

OpenStudy (anonymous):

thank you very much!

OpenStudy (aum):

You are welcome.